OpenStudy (anonymous):

Given the equation 4x + 7 = 3(2x − 5), solve for the variable. Explain each step and justify your process

OpenStudy (hybrik):

To do the first part, simplify the parathenses

OpenStudy (hybrik):

4x + 7 = 6x-15

OpenStudy (hybrik):

Then move the X values around: 4x-(6x)=-15-(7)

OpenStudy (anonymous):

What do you do next ?

OpenStudy (triciaal):

simplify what you have then divide each side by the coefficient of the variable to get one time the variable on one side x =

OpenStudy (triciaal):